# supplement

### v1

@@ 2.4 Optimization of hyper-parameters
In order to ensure the convergence rate of GL-ABC-MCMC, we need to optimize the hyperparameters (i.e. the global frequency $\gamma$, the batch size $N_{b}$). The expected squared jump distance (ESJD) (Pasarica and Gelman, 2010; Atchadé et al., 2011; Roberts and Rosenthal, 2014; Yang et al., 2020) is utilized to adjust the proposal distribution in the MCMC algorithm to improve the algorithm’s mixing and exploration. The ESJD in one dimension is defined as
@@
where $\rho_{1}$ is the first-order autocorrelation of the Markov chain. Maximizing the ESJD is equivalent to minimizing the first-order auto-correlation and thus maximizing the efficiency if the higher-order autocorrelations are monotonically increasing with respect to the first-order auto-correlation (Pasarica and Gelman, 2010). Roberts and Rosenthal (2014) demonstrated that maximizing the ESJD is equivalent to minimizing the asymptotic variance of diffusion limits of MCMC methods in certain scenarios. Yang et al. (2020) considered maximizing ESJD to optimal scaling of random-walk Metropolis algorithms on general target distributions, and showed that the asymptotically optimal acceptance rate $0.234$ can be obtained under general realistic sufficient conditions on the target distribution.
For high-dimensional problems, Pasarica and Gelman (2010) defined ESJD as
$\text{ESJD}=E[\|{\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t})\|_{\Sigma^{-1}}^{2}]=2\operatorname{tr}\left(I-\Sigma^{-1}E\left[({\boldsymbol{\theta}}_{t+1}-\bar{{\boldsymbol{\theta}}})({\boldsymbol{\theta}}_{t}-\bar{{\boldsymbol{\theta}}})^{\top}\right]\right),$ (2.3)
where $\bar{{\boldsymbol{\theta}}}$ and $\Sigma$ are the mean and covariance matrix of ${\boldsymbol{\theta}}$ from the stationary distribution. Maximizing (2.3) is equivalent to minimizing the first-order autocorrelation of the Markov chain. However, the covariance matrix $\Sigma$ is typically unknown until the target distribution is well-estimated. In this paper, we define ESJD based on the $D$-criterion as follows:
$\text{ESJD}=\left|\operatorname{E}\left[({\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t})({\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t})^{\top}\right]\right|^{\frac{1}{p}}=\left(|2\Sigma|\left|I-\Sigma^{-1}E\left[({\boldsymbol{\theta}}_{t+1}-\bar{{\boldsymbol{\theta}}})({\boldsymbol{\theta}}_{t}-\bar{{\boldsymbol{\theta}}})^{\top}\right]\right|\right)^{\frac{1}{p}},$ (2.4)
where $p$ is the dimension of ${\boldsymbol{\theta}}$. This accounts for different scales of the parameter dimensions without requiring the estimation of $\Sigma$. In extreme cases, Equation (2.4) takes the maximum value $|2\Sigma|^{1/p}$ when ${\boldsymbol{\theta}}_{t+1}$ and ${\boldsymbol{\theta}}_{t}$ are independent, and takes the minimum value 0 when ${\boldsymbol{\theta}}_{t+1}$ and ${\boldsymbol{\theta}}_{t}$ are equal. When $\Sigma$ is a diagonal matrix, Equation (2.4) measures the geometric mean of one-dimensional ESJD from different dimensions. Furthermore, considering the cost associated with different proposal mechanisms, we define a unit cost version of ESJD (cESJD) that accounts for computational cost as follows:
$\text{cESJD}=\frac{\text{ESJD}}{\mathcal{C}},$ (2.5)
@@
To enhance the performance of the GL-ABC-MCMC algorithm, we propose using a sequential optimization algorithm to select an optimal combination of hyper-parameters. This process involves selecting a set of candidate points in the hyper-parameter space based on specific rules, such as informative priors or non-informative uniform density. For each hyper-parameter combination, a short run of the GL-ABC-MCMC algorithm is performed, and the ESJD value is calculated. The best hyper-parameter combination, which yields the highest ESJD, is then used as the center of the parameter space, and the parameter space is narrowed down accordingly. This process is repeated until convergence is achieved or a predetermined requirement is met.
When there is a lack of prior information about hyper-parameters, utilizing uniform design (UD) (Fang et al., 2018) to explore the hyper-parameter space is robust and efficient. UD is constructed based on the number-theoretic method or quasi-Monte Carlo method and possesses good space-filling properties. Yang and Zhang (2021) proposed a method for hyper-parameter optimization via sequential uniform designs. By virtue of its space-filling property, uniform design ensures adequate exploration of the entire parameter space. Compared to random sampling, uniform design can achieve similar exploration effects with fewer sample points, thereby saving experimental resources and costs. The detailed algorithm is shown in the Supplement LABEL:supp:sec-sequential, Algorithm LABEL:alg:sequential.
2.5 Properties
@@
[…]
@@
A Markov kernel $Q$ with invariant probability measure $\pi$ is $V$-geometrically ergodic if there exit constants $\rho\in(0,1)$ and $M<\infty$ such that, for all $x\in\mathbb{X}$ and $k\in\mathbb{N}$, $\|Q^{k}(x,\cdot)-\pi\|_{V}\leq M\{V(x)+\pi(V)\}\rho^{k}$.
In the ABC-i-SIR algorithm, the importance proposal distribution is $\lambda({\boldsymbol{z}})\triangleq q({\boldsymbol{\theta}})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})$. The Markov chain generated by ABC-i-SIR admits the following Markov kernel:
@@
where $w({\boldsymbol{z}})=\pi({\boldsymbol{\theta}})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})/\lambda({\boldsymbol{z}})$ is the unnormalized importance weight function. We denote the function as $w({\boldsymbol{z}})=\pi_{\varepsilon}({\boldsymbol{z}})\cdot\lambda(w)/\lambda({\boldsymbol{z}})$, where $\lambda(w)$ is the normalizing constant of the distribution $\pi_{\varepsilon}({\boldsymbol{z}})$. Next, we present two propositions for the ABC-i-SIR algorithm.
@@
Proposition 1 shows that the ABC-i-SIR algorithm operates as a systematic-scan two-stage Gibbs sampler, with the marginal distribution of the target being $\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})$. Proposition 2 establishes that the MCMC kernel of ABC-i-SIR $P_{N_{b}}$ is $V$-geometrically ergodic. Here, the assumption $|w|_{\infty}<\infty$ implies that the importance proposal must cover the target distribution, that is, $q({\boldsymbol{\theta}})>0$ for all ${\boldsymbol{\theta}}\in\{{\boldsymbol{\theta}}:\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})\}$, which is a standard assumption for importance sampling. In practice, it suffices to set the importance proposal to cover the prior region. The proofs of Propositions 1 and 2 are shown in the Supplement LABEL:supp:sec-ABC-i_SIR and LABEL:supp:sec-Uniform_geometric_ergodicity_ABC-i-SIR.
@@
$\|K_{N_{b}}^{k}({\boldsymbol{z}},\cdot)-\pi_{\varepsilon}(\cdot)\|_{V}\leq c_{K_{N_{b}}}\{\pi_{\varepsilon}(V)+V({\boldsymbol{z}})\}\rho_{K_{N_{b}}}^{k},$ (2.6)
@@
Theorem 1 establishes that the kernel of GL-ABC-MCMC is $V$-uniformly geometrically ergodic under Assumptions 1 and 2. Specifically, Assumption 1(ii) requires that the Markov kernel $Q$ satisfies a Foster-Lyapunov drift condition for the function $V$, which is commonly met by traditional MCMC kernels like the Metropolis-Hastings kernel. Assumption 2 stipulates that the (normalized) importance weights are upper bounded on the sets $V_{r}$. This is considered as a mild condition: if $\mathbb{Z}=\mathbb{R}^{d}$ and $V$ is norm-like, then the level sets $V_{r}$ are compact, and $w(\cdot)$ is bounded on $V_{r}$ as long as the importance proposal covers the prior region. The condition $\operatorname{Var}_{\lambda}(w)/\{\lambda(w)\}^{2}\leq\infty$ implies that the variance of the importance weights is bounded, which corresponds to the $\chi^{2}$-distance between the proposal and target distributions, playing a crucial role in the performance of SIR methods. Notably, Assumptions 1 and 2 do not require the identification of small sets for the rejuvenation kernel $Q$, which is a necessary condition for the local kernel $Q$ to satisfy $V$-geometric ergodicity. Finally, Theorem 1 implies that the mixing rate $\rho_{K_{N_{b}}}^{k}$ of the GL-ABC-MCMC kernel $K_{N_{b}}$ exceeds that of the local kernel $Q$ in some sense.
3 Two adaptive schemes
In this section, we propose two adaptive schemes for ABC-i-SIR and local proposals to improve the inference of complex models.
3.1 An adaptive global proposal with normalizing flows
In Section 2.3, we employ the likelihood-free SIR technique iteratively to construct an efficient global kernel for ABC-MCMC. The proposal distribution plays a crucial role in importance sampling, influencing the efficiency of the sampling process. Essentially, the proposal determines where samples are drawn from, and its similarity to the target distribution significantly impacts the quality of estimates. An effective proposal reduces variability and yields more reliable results. The calculation of weights depends on how well the proposal aligns with the target, and optimal performance is achieved when the proposal closely approximates the target distribution. In this section, we enhance the importance proposal distribution of ABC-i-SIR to closely match the target distribution by training a normalizing flow throughout the sampling process.
Let $u\in\mathbb{R}^{d}$ be a random variable with a known and tractable probability density function $p_{B}$. Suppose $T$ is an invertible and differentiable function; then the density of $\nu=T(u)$ can be evaluated by
$p_{T}(\nu)=p_{B}(f(\nu))|\text{det}\nabla f(\nu)|=p_{B}(f(\nu))|\text{det}\nabla T(f(\nu))|^{-1},$ (3.7)
where $f$ is the inverse function of $T$ (i.e. $u=f(T(u))$), $\nabla f(\nu)=\partial f(\nu)/\partial\nu$ and $\nabla T(u)=\partial T(u)/\partial u$. Intuitively, if the transformation $T$ can be arbitrarily complex, we can generate any distribution $p_{T}$ from any base distribution $p_{B}$. A normalizing flow is a transformation of a simple distribution into a complex one by the composition of a sequence of invertible and differentiable functions. In order to generate our target distribution, we often parameterize a flow $T$ with parameter $\beta$, and obtain an optimal invertible map $T^{*}$ by minimizing the KL divergence between the ABC posterior $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})$, that is,
[…]
To estimate of gradient of $D_{\text{KL}}^{\text{F}}(\pi_{\varepsilon}\|p_{T})$, evaluating $p_{T}({\boldsymbol{\theta}})$ is straightforward. However, obtaining samples from $\pi_{\varepsilon}$ poses a challenge. Gabrié et al. (2022) suggest running $M$ parallel Markov chains, using their samples at each iteration as approximate samples from $\pi_{\varepsilon}$. This method, however, demands more computational resources than running a single chain. Here, we propose a novel approach that fully recycles all $N$ candidates sampled at each iteration of ABC-i-SIR. We treat the weighted particles $\left\{\theta_{t,i}^{*},w_{t,i}\right\}$ as the samples from $\pi_{\varepsilon}$. These particles $\left\{\theta_{t,i}^{*},w_{t,i}\right\}$ are then employed to approximate the integral in $D_{\text{KL}}^{\text{F}}(\pi_{\varepsilon}\|p_{T})$ (i.e., the loss function of $T$),
$L(T;\left\{\theta_{t,i}^{*},w_{t,i}\right\})=-\sum_{i=1}^{N}w_{i}\log(p_{T}({\boldsymbol{\theta}}_{(i)}^{*})).$
The constant $C$ in Equation (3.8) is omitted here since it does not affect the optimization of $T$. Algorithm 2 outlines a single stage of ABC-i-SIR algorithm enhanced with NF.
Algorithm 2 Single stage of ABC-i-SIR augmented with NF
Procedure ABC-i-SIR-NF($({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t}),q,P_{\boldsymbol{\theta}},K_{\varepsilon}$). Input: Previous state $({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$, normalizing flow $T$, simulator $P_{\boldsymbol{\theta}}$, kernel $K_{\varepsilon}$, prior $\pi({\boldsymbol{\theta}})$, step size of gradient descent $r$, number of steps to collect data $S$.
Output: New state $({\boldsymbol{\theta}}_{t+1},\boldsymbol{x}_{{t+1}})$, map $T$.
1:  ${\boldsymbol{\theta}}_{(0)}^{*}={\boldsymbol{\theta}}_{t}$, draw ${\boldsymbol{\theta}}_{(1:N_{b})}^{*}$ from $p_{T}(\cdot)$.
2:  Generate simulate data $\boldsymbol{x}_{(i)}^{*}$ from $P_{{\boldsymbol{\theta}}_{(i)}^{*}}$ for $i=0,\ldots,N_{b}$.
3:  Compute weight $w_{i}=w({\boldsymbol{\theta}}_{(i)}^{*})/\sum_{i=1}^{N_{b}}w({\boldsymbol{\theta}}_{(i)}^{*})$, for $i=0,\ldots,N_{b}$, where $w({\boldsymbol{\theta}}_{(i)}^{*})=\pi({\boldsymbol{\theta}}_{(i)}^{*})K_{\varepsilon}(\boldsymbol{x}_{(i)}^{*},\boldsymbol{y})/q({\boldsymbol{\theta}}_{{(i)}^{*}})$. Collect $\left\{\left({\boldsymbol{\theta}}_{(i)}^{*},w({\boldsymbol{\theta}}_{(i)}^{*})\right)\right\}_{i=1:N_{b}}$ into the train data set $\mathcal{D}$.
4:  Sample ${\boldsymbol{\theta}}_{t+1}$ from $\left\{{\boldsymbol{\theta}}_{(i)}^{*}\right\}_{i=0:N_{b}}$ with weight $\left\{w_{i}\right\}_{i=0:N_{b}}$.
5:  if $t\mod S$ = 0 then
6:     Normalize the weight in $\mathcal{D}$, $W_{i}=w({\boldsymbol{\theta}}_{(i)}^{*})/\sum_{i=1}^{S\cdot N_{b}}w({\boldsymbol{\theta}}_{(i)}^{*})$, for $i=1,\ldots,S\cdot N_{b}$.
7:     Compute $L(T)=-\sum_{i=1}^{S\cdot N_{b}}W_{i}\log(p_{T}({\boldsymbol{\theta}}_{(i)}^{*}))$, update the parameter $\beta$ of $T$, $\beta\leftarrow\beta-r\nabla_{\beta}L(T)$.
8:     Initialize $\mathcal{D}$ to an empty set.
9:  end if
3.2 A gradient based local proposal
In practice, it is common to employ a $p$-dimensional normal distribution as a proposal for local MCMC steps (e.g.$q({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})=\mathcal{N}({\boldsymbol{\theta}}^{*};{\boldsymbol{\theta}},\sigma^{2}\boldsymbol{I}_{p})$). The scale parameter of the normal distribution significantly influences the efficiency and performance of the algorithm, especially in high-dimensional parameter spaces. Designing an appropriate scale poses a challenge. It’s necessary to design a general proposal mechanism which provides large proposal transitions with a high probability of acceptance. Metropolis adjusted Langevin algorithm (MALA) (Grenander and Miller, 1994; Roberts and Tweedie, 1996; Roberts and Rosenthal, 1998; Girolami and Calderhead, 2011) combines Langevin dynamics with a Metropolis-Hastings correction to create an adaptive and efficient MCMC algorithm.
The Langevin diffusion process is defined by the stochastic differential equation (SDE):
$d{\boldsymbol{\theta}}(t)=\nabla_{\boldsymbol{\theta}}\mathcal{L}\{{\boldsymbol{\theta}}(t)\}dt/2+d\boldsymbol{W}(t),$ (3.9)
where $\mathcal{L}({\boldsymbol{\theta}})$ denotes the logarithm of target distribution, and $\boldsymbol{W}$ denotes a $p$-dimensional Brownian motion. The diffusion is irreducible, strong Feller and aperiodic, with stationary distribution $\pi_{T}(\cdot)\propto\exp(\mathcal{L}(\cdot))$. This process can be regarded as a continuous-time sampling method. Unfortunately, its implementation is infeasible in practice. The common method is to use discretized approximations via Euler-Maruyama discretization:
${\boldsymbol{\theta}}^{*}={\boldsymbol{\theta}}+\eta^{2}\nabla_{\boldsymbol{\theta}}\mathcal{L}({\boldsymbol{\theta}})/2+\eta z,$ (3.10)
where $\eta$ is the step-size and $z\sim\mathcal{N}(0,\boldsymbol{I}_{p})$. However, when discretizing the diffusion, some bias is introduced, and convergence to the invariant distribution $\pi_{T}$ is no longer guaranteed. MALA employs a Metropolis-Hastings rejection-acceptance step after every iteration to ensure the convergence property. The discrete form of Langevin diffusion serves as an instrumental proposal distribution. In particular, $q\left(\boldsymbol{\theta}^{*}\mid\boldsymbol{\theta}\right)=\mathcal{N}\left\{\boldsymbol{\theta}\mid{\boldsymbol{\theta}}+\eta^{2}\nabla_{\boldsymbol{\theta}}\mathcal{L}({\boldsymbol{\theta}})/2,\eta^{2}\boldsymbol{I}_{p}\right\}$ and the corresponding acceptance probability is $\alpha({\boldsymbol{\theta}},{\boldsymbol{\theta}}^{*})=\min\left\{1,\frac{\exp(\mathcal{L}({\boldsymbol{\theta}}^{*}))q({\boldsymbol{\theta}}\mid{\boldsymbol{\theta}}^{*})}{\exp(\mathcal{L}({\boldsymbol{\theta}}))q({\boldsymbol{\theta}}^{*}\mid{\boldsymbol{\theta}})}\right\}$.
[…]
$p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})=\sum_{s=1}^{S}K_{\varepsilon}(f({{\boldsymbol{\theta}}},\omega_{s}),\boldsymbol{y}),$ (3.11)
where $\omega_{s}$ ($s=1,\ldots,S$) are $S$ random seeds. Meeds et al. (2015) suggest approximating the conditional distribution $p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})$ by $K_{\varepsilon}(f({\boldsymbol{\theta}},\omega_{m}),\boldsymbol{y})$ under the assumption of a very small threshold, $\varepsilon$. Here, $\omega_{m}$ is the seed that ensures $K_{\varepsilon}(f({\boldsymbol{\theta}},\omega_{m}),\boldsymbol{y})$ is the largest among $\left\{K_{\varepsilon}(f({\boldsymbol{\theta}},\omega_{i}),\boldsymbol{y})\right\}_{i=1:S}$. The gradient of log-likelihood is then estimated by
$\nabla_{\boldsymbol{\theta}}\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))=\frac{\log(K_{\varepsilon}(f({{\boldsymbol{\theta}}}+d_{\boldsymbol{\theta}},\omega_{m}^{+}),\boldsymbol{y}))-\log(K_{\varepsilon}(f({{\boldsymbol{\theta}}}-d_{\boldsymbol{\theta}},\omega_{m}^{-}),\boldsymbol{y}))}{2d_{\boldsymbol{\theta}}},$ (3.12)
where $\omega_{m}^{+}=\arg\max_{\omega_{i}\in\{\omega_{1:S}\}}K_{\varepsilon}(f({{\boldsymbol{\theta}}}+d_{\boldsymbol{\theta}},\omega_{i}),\boldsymbol{y})$ and $\omega_{m}^{-}=\arg\max_{\omega_{i}\in\{\omega_{1:S}\}}K_{\varepsilon}(f({{\boldsymbol{\theta}}}-d_{\boldsymbol{\theta}},\omega_{i}),\boldsymbol{y})$. This method is denoted as $\text{CRN}_{max}$. In practice, achieving efficient ABC methods with a very small $\varepsilon$ is challenging. Here, we estimate the partial derivative with respect to ${\boldsymbol{\theta}}$ without assuming a small threshold as follows:
$\nabla_{\boldsymbol{\theta}}\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))=\frac{\log\left(\sum_{s=1}^{S}K_{\varepsilon}(f({{\boldsymbol{\theta}}}+d_{\boldsymbol{\theta}},\omega_{s}),\boldsymbol{y})\right)-\log\left(\sum_{s=1}^{S}K_{\varepsilon}(f({{\boldsymbol{\theta}}}-d_{\boldsymbol{\theta}},\omega_{s}),\boldsymbol{y})\right)}{2d_{\boldsymbol{\theta}}},$ (3.13)
where $d_{\boldsymbol{\theta}}$ is a small perturbation. The detailed finite difference stochastic approximation with CRN ($\text{CRN}_{mean}$) is shown in the Supplement LABEL:supp:sec-CRN.
However, in the tail region of the posterior distribution, the logarithmic function can amplify errors associated with Monte Carlo integration, leading to significantly inaccurate gradient estimates. To address this, we employ a Gaussian distribution to model the likelihood based on simulated data, achieving a more stable and accurate estimator when simulations are scarce for estimating the integral. By utilizing the Gaussian kernel as the kernel density function, we derive the gradient estimate as follows:
$\nabla_{\boldsymbol{\theta}}\log(p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}}))=\frac{1}{2d_{\boldsymbol{\theta}}}\left(-\frac{1}{2}\log\left(\frac{\hat{\sigma}_{{\boldsymbol{\theta}}^{+}}^{2}+\varepsilon^{2}}{\hat{\sigma}_{{\boldsymbol{\theta}}^{-}}^{2}+\varepsilon^{2}}\right)-\frac{(\boldsymbol{y}-\hat{\mu}_{{\boldsymbol{\theta}}^{+}})^{2}}{2(\hat{\sigma}_{{\boldsymbol{\theta}}^{+}}^{2}+\varepsilon^{2})}+\frac{(\boldsymbol{y}-\hat{\mu}_{{\boldsymbol{\theta}}^{-}})^{2}}{2(\hat{\sigma}_{{\boldsymbol{\theta}}^{-}}^{2}+\varepsilon^{2})}\right),$ (3.14)

### v2

@@ 3.2 Sequential optimization of hyper-parameters
In order to ensure the convergence rate of GL-ABC-MCMC, we need to optimize the hyperparameters (i.e. the global frequency $\gamma$, the batch size $N_{b}$). The expected squared jump distance (ESJD) (pasarica2010adaptivelyESJD; atchade2011towardsESJD; roberts2014minimisingESJD; yang2020optimalESJD) is utilized to adjust the proposal distribution in the MCMC algorithm to improve the algorithm’s mixing and exploration. The ESJD in one dimension is defined as
@@
where $\rho_{1}$ is the first-order auto-correlation of the Markov chain. Maximizing ESJD is equivalent to minimizing the first-order auto-correlation and thus maximizing the efficiency if the higher order auto-correlations are monotonically increasing with respect to the first-order auto-correlation (pasarica2010adaptivelyESJD). roberts2014minimisingESJD demonstrated that maximizing the ESJD is equivalent to minimizing the asymptotic variance of diffusion limits of MCMC methods in some certain scenarios.
yang2020optimalESJD consider maximizing ESJD to optimal scaling of random-walk Metropolis algorithms on general target distributions, and show the asymptotically optimal acceptance rate $0.234$ can be obtained under general realistic sufficient conditions on the target distribution.
For high-dimensional problems, pasarica2010adaptivelyESJD defined ESJD as
$\text{ESJD}=E[\|{\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t})\|_{\Sigma^{-1}}^{2}]=2\operatorname{tr}\left(I-\Sigma^{-1}E\left[({\boldsymbol{\theta}}_{t+1}-\bar{{\boldsymbol{\theta}}})({\boldsymbol{\theta}}_{t}-\bar{{\boldsymbol{\theta}}})^{\top}\right]\right),$ (9)
where $\bar{{\boldsymbol{\theta}}}$ and $\Sigma$ are the mean and covariance matrix of $\boldsymbol{\theta}$} from the stationary distribution. Maximizing (9) is equivalent to minimizing the first-order auto-correlation of the Markov chain. However, the covariance matrix $\Sigma$ is typically unknown until the target distribution is well-estimated. In this paper, we define ESJD based on the $D$-criterion as follows
$\text{ESJD}=\left|\operatorname{E}\left[({\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t})({\boldsymbol{\theta}}_{t+1}-{\boldsymbol{\theta}}_{t})^{\top}\right]\right|^{\frac{1}{p}}=\left(|2\Sigma|\left|I-\Sigma^{-1}E\left[({\boldsymbol{\theta}}_{t+1}-\bar{{\boldsymbol{\theta}}})({\boldsymbol{\theta}}_{t}-\bar{{\boldsymbol{\theta}}})^{\top}\right]\right|\right)^{\frac{1}{p}},$ (10)
where $p$ is the dimension of $\boldsymbol{\theta}$}. This accounts for different scales of the parameter dimensions while does not require the estimation of $\Sigma$. In extreme cases, Equation (10) takes the maximum value $|2\Sigma|^{1/p}$ when ${\boldsymbol{\theta}}_{t+1}$ and ${\boldsymbol{\theta}}_{t}$ are independent, and takes the minimum value 0 when ${\boldsymbol{\theta}}_{t+1}$ and ${\boldsymbol{\theta}}_{t}$ are equal. When $\Sigma$ is a diagonal matrix, Equation (10) measures the geometric mean of one dimension ESJD from different dimensions. Furthermore, considering the cost associated with different proposal mechanisms, we define a unit cost version of ESJD (cESJD) that takes the computational cost into account as:
$\text{cESJD}=\frac{\text{ESJD}}{\mathcal{C}},$ (11)
@@
To enhance the performance of the GL-ABC-MCMC algorithm, we propose using a sequential optimization algorithm to select an optimal hyper-parameter combination. This process involves selecting a set of candidate points in the hyper-parameter space based on specific rules, such as informative priors or non-informative uniform density. For each hyper-parameter combination, a short GL-ABC-MCMC algorithm is run, and the ESJD value is calculated. The best hyper-parameter combination, which yields the highest ESJD, is then used as the center of the parameter space, and the parameter space is narrowed down accordingly. This process is repeated until convergence is achieved or a predetermined requirement is met.
When there is a lack of prior information about hyper-parameters, using uniform design (UD) (fang1993number; fang2018UDbook) to explore the hyper-parameter space is robust and efficient. UD is constructed based on the number-theoretic method or quasi-Monte Carlo method, and it possesses good space-filling properties. yang2021hyperparameter proposed a method for hyper-parameter optimization via sequential uniform designs. By virtue of its space-filling property, uniform design ensures adequate exploration of the entire parameter space. Compared to random sampling, uniform design can achieve similar exploration effects with fewer sample points, thereby saving experimental resources and costs.
Algorithm 4 Sequential optimization algorithm
Input: Hyper-parameter space $\mathcal{K}$, maximum number of replications $R$, number of hyper-parameter combinations per stage $n_{s}$, iteration number of pilot GL-ABC-MCMC run $N_{p}$.
Output: The optimal hyper-parameter combination $\kappa^{*}$.
1: Initialize hyper-parameter set $\mathbb{K}=\emptyset$ and $\operatorname{ESJD}^{*}=0$.
2:for $r=1,\ldots,R$ do
3:  Select $\kappa_{1:n_{s}}$ in hyper-parameter space, $\kappa_{i}\notin\mathbb{K}$ for $i=1,\ldots,n_{s}$, and update $\mathbb{K}=\mathbb{K}\bigcup\kappa_{1:n_{s}}$.
4:  For each $\kappa_{i}$, run GL-MCMC with $N_{p}$ iterations and compute $\operatorname{ESJD}_{1:n_{s}}$.
5:  If $\operatorname{ESJD}_{m}>\operatorname{ESJD}^{*}$, update $\kappa^{*}=\kappa_{m}$ and $\operatorname{ESJD}^{*}=\operatorname{ESJD}_{m}$, where $m=\arg\max_{i}\{\operatorname{ESJD}_{1:n_{s}}\}$.
6:end for
The detailed algorithm is shown in Algorithm 4. The algorithm achieves adaptive exploration of the hyper-parameter space by sequentially adjusting the search space. At each stage, the search space is automatically adjusted to focus on the most promising region.
4 Properties
@@
Notations: Denote $\{$${\mathbb{N}^{*}}=\mathbb{N}\backslash\{0\}$. For a measurable function $f:\mathbb{X}\mapsto\mathbb{R}$, we define $|f|_{\infty}=$ $\sup_{x\in\mathbb{X}}|f(x)|$ and $\pi(f):=\int_{\mathbb{X}}f(x)\pi({d}x)$. For a function $V:\mathbb{X}\mapsto[1,\infty)$, we introduce the $V$-norm of two probability measures $\xi$ and $\xi^{\prime}$ on $(\mathbb{X},\mathcal{X}),\left\|\xi-\xi^{\prime}\right\|_{V}:=\sup_{|f(x)|\leq V(x)}\left|\xi(f)-\xi^{\prime}(f)\right|$. If $V\equiv 1,\|\cdot\|_{1}$ is equal to the total variation distance (denoted $\|\cdot\|_{{TV}}$). For simplicity, we denote the parameter-data pair $({\boldsymbol{\theta}},\boldsymbol{x})$ as $\boldsymbol{z}$}, ${\boldsymbol{z}}\in\mathbb{Z}=\Theta\times\mathbb{Y}$, and the ABC posterior distribution $\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})$ as $\pi_{\varepsilon}({\boldsymbol{z}}\mid\boldsymbol{y})$, which is defined on the measurable space $(\mathbb{Z},\mathcal{Z})$.
@@
A Markov kernel $Q$ with invariant probability measure $\pi$ is $V$-geometrically ergodic if there exist constants $\rho\in(0,1)$ and $M<\infty$ such that, for all $x\in\mathbb{X}$ and $k\in\mathbb{N}$, $\|Q^{k}(x,\cdot)-\pi\|_{V}\leq M\{V(x)+\pi(V)\}\rho^{k}$.
In the ABC-i-SIR algorithm, we extend the proposal distribution from $q({\boldsymbol{\theta}})$ to $\lambda({\boldsymbol{z}})\triangleq q({\boldsymbol{\theta}})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})$ by introducing the auxiliary variable $\boldsymbol{x}$, which represents simulated data generated under the parameter $\boldsymbol{\theta}$}. The Markov chain generated by ABC-i-SIR admits the following Markov kernel
@@
where $w({\boldsymbol{z}})=\pi({\boldsymbol{\theta}})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})/\lambda({\boldsymbol{z}})$ is the unnormalized importance weight function. We denote the function as $w({\boldsymbol{z}})=\pi_{\varepsilon}({\boldsymbol{z}}\mid\boldsymbol{y})\cdot\lambda(w)/\lambda({\boldsymbol{z}})$, where $\lambda(w)$ is the normalizing constant of the distribution $\pi_{\varepsilon}({\boldsymbol{z}}\mid\boldsymbol{y})$. Following the work of samsonov2022local-global, we demonstrate that the likelihood-free ABC-i-SIR algorithm can also be interpreted as a systematic-scan two-stage Gibbs sampler, which possesses uniform geometric ergodicity. The formal statements of these results are provided in Proposition 1 and Proposition 2, and the proofs are shown in the Appendix.
@@
In Proposition 2, the assumption $|w|_{\infty}<\infty$ implies that the importance proposal must cover the target distribution; that is, $q({\boldsymbol{\theta}})>0$ for all ${\boldsymbol{\theta}}\in\{{\boldsymbol{\theta}}:\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})>0\}$, which is a standard assumption for importance sampling. In practice, it suffices to set the importance proposal to cover the prior region.
It is well established that if both the global and local Markov kernels are geometrically ergodic, then their combination within the Global-Local ABC-MCMC framework inherits geometric ergodicity. In Proposition 2, we have shown that the Markov kernel underlying the ABC-i-SIR algorithm exhibits uniform geometric ergodicity. Building on this, Theorem 1 further establishes that when the ABC-i-SIR algorithm is used as the global Markov kernel, even if the local kernel does not satisfy $V$-geometric ergodicity, the resulting Global-Local ABC-MCMC kernel still attains $V$-geometric ergodicity. In other words, the mixing rate of the Global-Local ABC-MCMC kernel surpasses that of the local kernel alone, indicating improved convergence behavior.
@@
$\|K_{N_{b}}^{k}({\boldsymbol{z}},\cdot)-\pi_{\varepsilon}(\cdot)\|_{V}\leq c_{K_{N_{b}}}\{\pi_{\varepsilon}(V)+V({\boldsymbol{z}})\}\rho_{K_{N_{b}}}^{k},$ (12)
@@
It is important to emphasize that Assumption 1(ii) requires the local Markov kernel $Q$ to satisfy a Foster-Lyapunov drift condition with respect to some function $V$. This condition is commonly met by classical MCMC kernels such as the Metropolis-Hastings algorithm. However, the Foster-Lyapunov drift condition alone does not guarantee the geometric ergodicity of the local kernel. Additionally, Assumption 2 imposes an upper bound on the (normalized) importance weights within the level sets $V_{r}$. This is a mild and practical condition: for example, when $\mathbb{Z}=\mathbb{R}^{d}$ and $V$ is norm-like, the sets $V_{r}$ are compact, and the importance weights $w(\cdot)$ remain bounded as long as the importance proposal sufficiently covers the prior support. Moreover, the condition $\operatorname{Var}_{\lambda}(w)/\{\lambda(w)\}^{2}\leq\infty$ ensures bounded variance of the importance weights, which corresponds to a finite $\chi^{2}$-distance between the proposal and target distributions, playing a critical role in the performance of Sampling Importance Resampling (SIR) methods.
From a practical perspective, the global ABC-i-SIR kernel facilitates strong global moves by enabling efficient jumps across distant regions of the state space, while the local kernel focuses on detailed local exploration. Thus, even if the local kernel alone lacks $V$-geometric ergodicity, the combined Global-Local kernel can achieve overall $V$-geometric ergodicity, thereby enhancing the efficiency and robustness of the ABC-MCMC algorithm.